Diffusion Morning Report 10/05/06

Ted Berg tberg at stanford.edu
Thu Oct 5 06:20:34 PDT 2006


Good Morning All,
    Looking OK this morning. We(Mike) are pulling the Epi chamber today 
after much testing to hopefully solve base pressure issue. Injectors 
were changed on furnaces tylan1 and 3 yesterday. TC sheath changed on 
Thermco poly yesterday. Everything else looks ok for now. no progress on 
AMAT Epi.

> Diffusion Mike,Ray, Ted 	  	  	  	  	 
>   	Date: 	10/5/06 	THICKNESS-Change at 	  	 
> Tube Name/Number 	Status 	Comments 	  	  	 
> Tylan 1 	Up 	  	  	  	 
> Tylan 2 	Up 	  	  	  	 
> Tylan 3 	Up 	  	  	  	 
> Tylan 4 	Up 	  	  	  	 
> Tylan 5-BBr3 	Up 	  	  	  	 
> Tylan 6-POCl3 	Up 	  	  	  	 
> Tylan 7-TEOS2 	Up 	  	8.41/30 	  	 
> Tylan 8-SiGe 	Up 	  	6.72/30 	  	 
> Tylan 9-Poly 	Up 	  	.35/30 	  	 
> Tylan 10-Nitride 	Up 	  	5.45/30 	  	 
> Tylan 11-BPSG-LTO 	Up 	  	25.92/30 	  	 
> Tylan 12-FGA 	Up 	  	  	  	 
>   	  	  	  	  	 
> Thermco 1 	Up 	  	  	  	 
> Thermco 2 	Up 	  	  	  	 
> Thermco Poly 	Up 	  	  	  	 
> Thermco Nitride 	Up 	  	  	  	 
> Epi 	Down 	high base pressure 	  	  	 
>   	
> 	
> 	
> 	
> 	
> Burn Box 	up 	  	  	  	 
> ZYGO 	Up 	  	  	  	 
>   	up 	  	   	  	 
> GAAS 22 	  	  	  	  	 
> GAAS 23 	UP 	  	  	  	 
> SiH4 for furnaces 	up 	  	  	  	 
>
> 	  	B on line 320 psi 	 changeover working well  	  	 
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061005/31163d38/attachment.html>


More information about the diffusion mailing list